# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

### v2

@@ III X17 couplings to charm and strange quarks
The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=6.4\times 10^{-3}$ and $\left|\varepsilon_{s}\right|=2.0\times 10^{-3}$ while there are two best-fit solutions for $\left|\varepsilon_{u}\right|$, which are $6.0\times 10^{-2}$ and $6.5\times 10^{-2}$, respectively. We have $\chi^{2}_{\rm min}=0.46$ for the fitting. The fitting results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. Up to the $3\sigma$ range, the value of $\left|\varepsilon_{u}\right|$ is of the order $10^{-2}$, which is notably larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ Denton:2023gat . In other words, the value of $\left|\varepsilon_{u}\right|$ extracted from the decays of $D^{*}$, $\psi(2S)$, and $\phi$ mesons is in a serious tension with that extracted by ATOMKI measurements. The significant increase of $\left|\varepsilon_{u}\right|$ in this case can be attributed to the anomalously large branching ratio of $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ measured by by BESIII BESIII:2021vyq as illustrated by Fig. 2 and the right panel of Fig. 3. We remark that only the decay of $D^{*0}$ among our considered decay modes involve X17 couplings to the first generation of quarks.
IV Conclusions
@@
The confirmation of couplings between X17 boson and quarks remains a topic of interest to date. The hypothesis of X17 boson has continuously received attentions by the data of anomalous 8Be, 4He, and 12C decays over time. The accumulation of data has shifted the most favored coupling strengths from $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ Feng:2016jff to $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ without isospin effects, and $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ with both isospin mixing and breaking effects considered Denton:2023gat . The persistent decay anomalies outlined above have motivated the testing of X17 hypothesis in the decays of heavy-flavor mesons such as $D^{*}$ and $B^{*}$ mesons with the assumption $\varepsilon_{b}=\varepsilon_{s}=\varepsilon_{d}$ and $\varepsilon_{c}=\varepsilon_{u}$ Castro:2021gdf . We have revisited this study by focusing on the comparisons of theoretical predictions (with X17 boson mediated effects included) on decay ratios $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ with experimental data CLEO:2011mla ; BESIII:2021vyq . For the case of $R_{D_{s}^{*+}}$, we have found that only the updated values of $\varepsilon_{u,d}$ are consistent with the measurement while the earlier extracted parameter values $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ predict a $R_{D_{s}^{*+}}$ in tension with the measurement (see Fig. 1). Our finding corrects an error by Ref. Castro:2021gdf which concluded that the prediction by $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ is consistent with the measurement.
@@
To accommodate the measured $R_{D^{*0}}$, one requires a much larger $\left|\varepsilon_{u}\right|$ under the assumption $\varepsilon_{c}=\varepsilon_{u}$. On the other hand, raising the magnitude of $\varepsilon_{c}$ would affect $R_{D_{s}^{*+}}$ unless the magnitude of $\varepsilon_{s}$ is increased as well since $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ as indicated by Fig. 3 (left panel). Since $\varepsilon_{s}$ already departs from $\varepsilon_{d}$ in this case, we then also treat $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent parameters in our analysis. To determine the favored ranges for $\varepsilon_{u}$, $\varepsilon_{c}$ and $\varepsilon_{s}$, we have performed a fitting to the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}$, and $R_{\phi}$. The results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=6.4\times 10^{-3}$, $\left|\varepsilon_{s}\right|=2.0\times 10^{-3}$, and two solutions for $\left|\varepsilon_{u}\right|$, which are $6.0\times 10^{-2}$ and $6.5\times 10^{-2}$, respectively.
@@
It is noteworthy that the value of $\left|\varepsilon_{u}\right|$ remains of the order $10^{-2}$ up to the $3\sigma$ range. Remarkably, this value is much larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|=(0.5-0.9)\times 10^{-3}$ Denton:2023gat . Such a tension is caused by the large deviation of the measured $R_{D^{*0}}$ from the Standard Model prediction as shown in Table 2. This is also reflected by the right panel of Fig. 3, where both green bands significantly deviate from the line $\varepsilon_{u}=0$. In conclusion, we have studied effects of the proposed X17 boson to the decays of $D^{*0}$, $D^{*+}_{s}$, $\psi(2S)$, and $\phi$ mesons, which involve X17 boson couplings to both the first and second generations of quarks. The above-mentioned tension on the value of $\varepsilon_{u}$ deserves further scrutiny. We note that the decay $D^{*\pm}\to D^{\pm}e^{+}e^{-}$ remains to be measured. Furthermore the huge deviation of the measured $D^{*0}\to D^{0}e^{+}e^{-}$ branching ratio to its Standard Model predicted value awaits confirmations from updated measurements.

### v3

@@ III X17 couplings to charm and strange quarks
Since more statistics for $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ is required for deciphering the nature of large excess in $R_{D^{*0}}$, it is of interest to perform two fittings with one that includes only $D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-}$, $\psi(2S)\to\eta_{c}e^{+}e^{-}$, and $\phi\to\eta e^{+}e^{-}$ data and the other one that includes $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ data as well. The first fitting only involves $\varepsilon_{c}$ and $\varepsilon_{s}$ which are not constrained by ATOMKI measurements. The result of the first fitting is shown in Fig. 5. We employ the $\chi^{2}$ statistical test for fitting the coupling parameters $\varepsilon_{c}$, and $\varepsilon_{s}$ with the three relevant measurements given in Table 2. For $\phi\to\eta e^{+}e^{-}$ decay, we only employ the data by KLOE-2 KLOE-2:2014hfk since its uncertainty is much smaller than those of earlier measurements. We minimize the $\chi^{2}$ defined as
$\chi^{2}=\sum_{i=1}^{3}\frac{\left(R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\lambda\right)-R_{i}^{\rm ob}\right)^{2}}{\sigma_{i}^{2}}+\frac{\left(\lambda-\lambda_{\rm best\,fit}\right)^{2}}{\sigma_{\lambda}^{2}},$ (24)
where $R_{i}^{\rm th}$ denotes the theoretically predicted value for the $i$-th decay ratio with $i=D^{*+}_{s},\ \psi(2S),\ {\rm and}\ \phi$. These ratios depend on the parameters $\varepsilon_{c}$, $\varepsilon_{s}$, and $\lambda$. $R_{i}^{\rm ob}$ refers to the observed value for the $i$-th decay ratio with $\sigma_{i}$ the corresponding uncertainty for the measurement. The parameter $\lambda$ is defined in Eq. (11) with its range given by Eq. (13). We hence take $\lambda_{\rm best\,fit}=-0.289$ GeV-1 and $\sigma_{\lambda}=0.016$ GeV-1. Our fitting uses the measured decay ratios of $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons as outlined in Table 2. We have $\chi^{2}_{\rm min}=0.45$ for the fitting.
[…]
The best-fit values are given by $\varepsilon_{c}=\pm 7.6\times 10^{-3}$ and $\varepsilon_{s}=\mp 2.4\times 10^{-3}$. We note that the best-fit value for $\varepsilon_{c}$ is significantly larger in magnitude than that of $\varepsilon_{u}$ obtained from fitting the ATOMKI data Denton:2023gat (see also the caption of Table 1). On the other hand, the allowed range for $\varepsilon_{c}$ is still too large to indicate whether these new physics couplings are universal in quark generations or not.
Let us now include $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ into the fitting. We note that only this decay among our considered modes involve X17 coupling to the first generation of quarks, i.e., $\varepsilon_{u}$. The fitting results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 6. The best-fit values for the coupling parameters are $\left|\varepsilon_{c}\right|=7.6\times 10^{-3}$ and $\left|\varepsilon_{s}\right|=2.4\times 10^{-3}$ while there are two best-fit solutions for $\left|\varepsilon_{u}\right|$, which are $6.1\times 10^{-2}$ and $6.7\times 10^{-2}$, respectively. To understand the origin of double solutions for $\varepsilon_{u}$, let us apply the narrow width approximation, Eq. (8), to Eq. (7) and write the theoretical prediction to $D^{*0}(c\bar{u})$ decay ratio as
$R^{\rm th}_{D^{*0}}=R^{\gamma}_{D^{*0}}+\kappa\,(A\,\varepsilon_{c}+B\,\varepsilon_{u})^{2},$ (25)
here $A\equiv 1/m_{D^{*0}}$ and $B\equiv 1/m_{u}(m_{X}^{2})$ with $m_{u}(q^{2})^{-1}$ given by Eq. (10). The coefficient $\kappa\equiv K/D^{2}$ collects the kinematic factor $K\equiv\lambda^{3/2}(m_{D^{*0}}^{2},m_{D^{0}}^{2},m_{X}^{2})/(m_{D^{*0}}^{2}-m_{D^{0}}^{2})^{3}$ and the photon-contribution normalization $D\equiv e_{c}/m_{D^{*0}}+e_{u}/m_{u}(0)$ from Eqs. (7) and (9). For a given value for $\chi^{2}_{D^{*0}}(\varepsilon_{u},\varepsilon_{c})\equiv\big(R_{D^{*0}}^{\text{th}}(\varepsilon_{u},\varepsilon_{c})-R_{D^{*0}}^{\exp}\big)^{2}/\sigma_{D^{*0}}^{2}$, we apply Eq. (25) and obtain $A\varepsilon_{c}+B\varepsilon_{u}=\pm\sqrt{C}$ where $C>0$ is a constant determined by the given $\chi^{2}_{D^{*0}}$ value and the uncertainty $\sigma_{D^{*0}}^{2}$ in the measurement. We then obtain the solutions for $\varepsilon_{u}$ as $\varepsilon_{u}=(-A\varepsilon_{c}\pm\sqrt{C})/B$. That $\pm$” leads to the existence of double solutions for $\varepsilon_{u}$, i.e., the fitting returns two values for $|\varepsilon_{u}|$ (here $|\varepsilon_{u}|\simeq 0.061$ and $|\varepsilon_{u}|\simeq 0.067$) while keeping $|\varepsilon_{c}|\simeq 7.6\times 10^{-3}$ and $|\varepsilon_{s}|\simeq 2.4\times 10^{-3}$.
Table 3: Comparison of best-fit parameters with and without including the $R_{D^{*0}}$ data
Scenario $\varepsilon_{c}$ $\varepsilon_{s}$ $\varepsilon_{u}$ Comment
Fit without $R_{D^{*0}}$ $\pm 7.6\times 10^{-3}$ $\mp 2.4\times 10^{-3}$
Fit with $R_{D^{*0}}$ $\pm 7.6\times 10^{-3}$ $\mp 2.4\times 10^{-3}$ $\pm 6.1\times 10^{-2}$$\mp 6.7\times 10^{-2}$ tension with ATOMKI
The fitted $\varepsilon_{c}$ and $\varepsilon_{s}$ are the same as those in the previous fitting with $R_{D^{*0}}$ excluded. Such a result can be understood from Fig. 3. On the left panel of the figure, the narrow bands resulting from CLEO CLEO:2011mla and KLOE-2 KLOE-2:2014hfk experiments essentially determine the favored ranges for $\varepsilon_{c}$ and $\varepsilon_{s}$. When $\varepsilon_{c}$ is frozen to a small range as the above, the favored ranges for $\varepsilon_{u}$ can be seen from the right panel of Fig. 3 as the intersections of the tiny $\varepsilon_{c}$ favored range with the green bands dictated by the measurement on $R_{D^{*0}}$ BESIII:2021vyq . Clearly, $R_{D^{*0}}$ data does not affect the favored range for $(\varepsilon_{c},\varepsilon_{s})$ unless the above intersections do not occur. Up to the $3\sigma$ range, the value of $\left|\varepsilon_{u}\right|$ is of the order $10^{-2}$, which is notably larger than that determined from the anomalous decays of $^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and $^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ Denton:2023gat . In other words, the value of $\left|\varepsilon_{u}\right|$ extracted from the decays of $D^{*0}$, $D^{*+}_{s}$, $\psi(2S)$, and $\phi$ mesons is in a serious tension with that extracted by ATOMKI measurements. The significant increase of $\left|\varepsilon_{u}\right|$ in this case can be attributed to the anomalously large branching ratio of $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ measured by BESIII BESIII:2021vyq as illustrated by Fig. 2 and the right panel of Fig. 3. We point out that both $\varepsilon_{u}$ branches in Eq. (25) give rise to the same $R_{D^{*0}}^{\text{th}}$. Hence, both produce the same $\chi^{2}_{D^{*0}}(\varepsilon_{u},\varepsilon_{c})$. Since none of $R_{D^{*+}_{s}}$, $R_{\psi(2S)}$, and $R_{\phi}$ depends on $\varepsilon_{u}$, therefore both $\varepsilon_{u}$ branches also lead to the same overall $\chi^{2}$. In other words, both solutions for $\varepsilon_{u}$ are physically acceptable. As a summary, we list the best-fit parameters obtained with and without the $R_{D^{*0}}$ data in Table 3.
Figure 7: The theoretically predicted values for $R_{D^{*+}}$ derived from Eqs. (6) and  (7) are shown. The scenarios RDa1 and RDa2 assume $\varepsilon_{c}=\varepsilon_{u}$, where RDa1 is calculated using $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$, and RDa2 is obtained by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ Denton:2023gat . The labels RDb1, RDb2, RDb3, and RDb4 represent the 1-sigma range of $\varepsilon_{c}$ from Fig. 5, with different assumptions about $\varepsilon_{d}$. For $\varepsilon_{d}=\mp 2.9\times 10^{-3}$, RDb1 corresponds to the same-sign scenario of $(\varepsilon_{c},\varepsilon_{d})$, while RDb2 represents the opposite-sign scenario. Similarly, for $\varepsilon_{d}=\mp 2.5\times 10^{-3}$, RDb3 and RDb4 correspond to the same-sign and opposite-sign scenarios, respectively. The SM prediction is denoted as SM_Photon.
With the above information on X17 coupling parameters, we give our predictions for the not-yet measured ratio $R_{D^{*+}}\equiv\Gamma(D^{*+}\to D^{+}e^{+}e^{-})/\Gamma(D^{*+}\to D^{+}\gamma)$ in Fig. 7. The SM_Photon is the Standard Model baseline with only the virtual-photon contribution in $\Gamma(D^{*+}\to D^{+}e^{+}e^{-})$. RDa1 and RDa2 give predictions assuming the generation-universality $\varepsilon_{c}=\varepsilon_{u}$. The former takes $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ while the latter adopts $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ Denton:2023gat . Both scenarios predict $R_{D^{*+}}$ above the SM value due to the additional X17 mediated amplitude. The vertical bands RDb1–RDb4 relax generation universality by scanning the $1\sigma$ range of $\varepsilon_{c}$ obtained from the combined fitting to $R_{D_{s}^{*+}}$, $R_{\phi(2S)}$, and $R_{\phi}$ given in Table 2 with $|\varepsilon_{d}|$ fixed at either $2.9\times 10^{-3}$ or $2.5\times 10^{-3}$ and different relative signs between $\varepsilon_{c}$ and $\varepsilon_{d}$ considered. The central value of each $R_{D^{*+}}$ prediction primarily depends on the magnitude of $\varepsilon_{c}$, i.e., when $|\varepsilon_{c}|$ is larger, the X17 contribution to $\Gamma(D^{*+}\to D^{+}e^{+}e^{-})$ is larger such that $R_{D^{*+}}$ is shifted upward; the opposite holds for a smaller $|\varepsilon_{c}|$. Meanwhile, the range of each RDb band is controlled by how $\varepsilon_{c}$ and $\varepsilon_{d}$ interferes each other. If $\varepsilon_{c}$ and $\varepsilon_{d}$ have the same sign, their associated amplitudes add up; if they have opposite signs, their corresponding amplitudes partially cancel each other. If a future measurement finds $R_{D^{*+}}$ significantly above the SM_Photon baseline and within the RDb bands as shown in Fig. 7, then the ATOMKI-sized $|\varepsilon_{c}|=|\varepsilon_{u}|\simeq(0.5–0.9)\times 10^{-3}$ would generally be too small to account for the observed enhancement. This then points toward the generation non-universality, i.e., $\varepsilon_{c}\neq\varepsilon_{u}$ and possibly $\varepsilon_{s}\neq\varepsilon_{d}$.
IV Summaries and Conclusions
@@
The confirmation of couplings between X17 boson and quarks remains a topic of interest to date. The hypothesis of X17 boson has continuously received attentions by the data of anomalous 8Be, 4He, and 12C decays over time. The accumulation of data has shifted the most favored coupling strengths from $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ Feng:2016jff to $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ without isospin effects, and $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ with both isospin mixing and breaking effects considered Denton:2023gat . The persistent decay anomalies outlined above have motivated the testing of X17 hypothesis in the decays of heavy-flavor mesons such as $D^{*}$ and $B^{*}$ mesons with the assumption $\varepsilon_{b}=\varepsilon_{s}=\varepsilon_{d}$ and $\varepsilon_{c}=\varepsilon_{u}$ Castro:2021gdf . We have revisited this study by focusing on the comparisons of theoretical predictions (with X17 boson mediated effects included) on decay ratios $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ with experimental data CLEO:2011mla ; BESIII:2021vyq . For the case of $R_{D_{s}^{*+}}$, we have found that only the updated values of $\varepsilon_{u,d}$ are consistent with the measurement while the earlier extracted parameter values $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ predict a $R_{D_{s}^{*+}}$ in a slight tension with the measurement (see Fig. 1). Our finding corrects an error in Ref. Castro:2021gdf which concluded that the prediction by $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ is consistent with the measurement.
@@
To accommodate the measured $R_{D^{*0}}$, one requires a much larger $\left|\varepsilon_{u}\right|$ under the assumption $\varepsilon_{c}=\varepsilon_{u}$. On the other hand, raising the magnitude of $\varepsilon_{c}$ would affect $R_{D_{s}^{*+}}$ unless the magnitude of $\varepsilon_{s}$ is increased as well since $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ with two parameters in opposite signs, as indicated by Fig. 3 (left panel). Since $\varepsilon_{s}$ already departs from $\varepsilon_{d}$ in this case, we then also treat $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent parameters in our analysis.
We have predicted the spectrum ${\rm d}R_{D^{*0}}/{\rm d}q^{2}$ assuming that the excess in $R_{D^{*0}}$ is due to the X17 boson. Our prediction is shown in Fig. 4 where the $D^{*0}\to D^{0}X\to D^{0}e^{+}e^{-}$ decay peak shows up in the second bin of the spectrum. Since the current statistics of Ref. BESIII:2021vyq is not yet sufficient to confirm or disconfirm this peak, we have fitted the X17 coupling parameters in two scenarios: one that considers only $R_{D_{s}^{*+}}$, $R_{\psi(2S)}$, and $R_{\phi}$ data, while the other includes the data of $R_{D^{*0}}$ as well. For the former scenario, the fitting result projected onto $(\varepsilon_{c},\varepsilon_{s})$ plane is shown in Fig. 5. The best-fit values for the coupling parameters are $\left|\varepsilon_{c}\right|=7.6\times 10^{-3}$, $\left|\varepsilon_{s}\right|=2.4\times 10^{-3}$. Although the best-fit value for $\varepsilon_{c}$ is significantly larger in magnitude than that of $\varepsilon_{u}$ obtained from fitting the ATOMKI data Denton:2023gat , the allowed range for $\varepsilon_{c}$ is still too large to verify whether these new physics couplings are independent of quark generations or not. For the latter scenario, the fitting also yields the range for $\varepsilon_{u}$ as given in Fig. 6. There are two solutions which correspond to the best-fit values of $|\varepsilon_{u}|$ as $6.1\times 10^{-2}$ and $6.7\times 10^{-2}$, respectively. The corresponding best-fit couplings, obtained with and without $R_{D^{*0}}$, are summarized in Table 3.
@@
It is noteworthy that the value of $\left|\varepsilon_{u}\right|$ shown in Fig. 6 remains of the order $10^{-2}$ up to the $3\sigma$ range. Remarkably, this value is much larger than that determined from the anomalous decays of $^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and $^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|=(0.5-0.9)\times 10^{-3}$ Denton:2023gat . Such a tension is caused by the large deviation of the measured $R_{D^{*0}}$ from the SM prediction as shown in Table 2. This is also reflected by the right panel of Fig. 3, where both green bands significantly deviate from the line $\varepsilon_{u}=0$.
We note that the decay channel $D^{*+}\rightarrow D^{+}e^{+}e^{-}$ has not been measured. We have predicted $R_{D^{*+}}$ based on our understanding of X17 related couplings $\varepsilon_{d}$ and $\varepsilon_{c}$ as given in Fig. 7. The central value and the uncertainty of $R_{D^{+*}}$ are both small by assuming the generation universality $\varepsilon_{c}=\varepsilon_{u}$. On the other hand, both the central values and the uncertainties of $R_{D^{+*}}$ become larger if we take the values of $\varepsilon_{c}$ as those given by the fitting range in Fig. 5.
In conclusion, we have studied the effects of the proposed X17 boson to the decays of $D^{*0}$, $D^{*+}_{s}$, $\psi(2S)$, and $\phi$ mesons, which involve X17 boson couplings to both the first and second generations of quarks. The above-mentioned tension on the value of $\varepsilon_{u}$ deserves further scrutiny., i.e., the huge deviation of the measured $D^{*0}\to D^{0}e^{+}e^{-}$ branching ratio to its SM predicted value awaits confirmations from updated measurements.